# Effective Polaron Dynamics of an Impurity Particle Interacting with a Fermi Gas

### v1

@@ 3.2 Effective coherent state
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ \|(\mathcal{N}+3)e^{aB}\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|$
$\leq C\|\eta_{s}\|\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|a}\|(\mathcal{N}+5)^{2}e^{\sigma B}\psi\|$
$\leq C\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma(e^{C\|\eta_{s}\|(1-\sigma)}-1)e^{\tilde{C}\|\eta_{s}\|\sigma}\|(\mathcal{N}+8)^{2}\psi\|$
$\leq C\|\eta_{s}\|^{-1}\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|e^{C\|\eta_{s}\|(1-\tau)}\|(\mathcal{N}+8)^{2}\psi\|$ (5.50)
where we used Cauchy-Schwarz with (5.16) and Lemma 5.1 in the first inequality, Proposition 2.6 and Lemma A.3 in the second inequality. Therefore it follows
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,2}\psi\|^{2}}$
$\leq CM^{2}N^{-\frac{4}{3}+2\delta}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+8)^{2}\psi\|$ (5.51)
@@
$\|I_{4,3}\psi\|\leq C\|\eta_{s}\|^{-1}MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|(e^{C\|\eta_{s}\|}-e^{C\|\eta_{s}\|\tau})\|(\mathcal{N}+3)\psi\|$ (5.52)
@@
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,3}\psi\|^{2}}\leq CMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+3)\psi\|.$ (5.53)
@@
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \sigma\|\mathcal{N}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\sigma\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)e^{\sigma B}\psi\|$
$\leq C(\|\eta_{s}\|+\|\eta_{s}\|^{2})MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \sigma\|(\mathcal{N}+1)^{2}e^{\sigma B}\psi\|$
$\leq C(\|\eta_{s}\|^{-1}+1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\ \left(e^{C\|\eta_{s}\|}+(1-C\|\eta_{s}\|\tau)e^{C\|\eta_{s}\|\tau}\right)\|(\mathcal{N}+3)^{2}\psi\|$
$\leq C(\|\eta_{s}\|^{-1}+1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\ e^{C\|\eta_{s}\|}\|(\mathcal{N}+3)^{2}\psi\|$ (5.54)
@@
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,4}\psi\|^{2}}$
@@
$\|I_{4,5}\psi\|$ $=\|2i\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\int_{0}^{\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{k,\alpha}e^{aB}\psi\|$
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ \|\mathcal{N}e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{k,\alpha}e^{aB}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|(1-a)}\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|(\mathcal{N}+3)\mathcal{R}^{\sigma-a}_{\alpha}(k)e^{aB}\psi\|^{2}\bigg{)}^{1/2}$
$\leq C\|\eta_{s}\|M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|(1-a)}(e^{C\|\eta_{s}\|(\sigma-a)}-1)\|(\mathcal{N}+8)^{2}e^{aB}\psi\|$
$\leq C\|\eta_{s}\|M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|}e^{C^{\prime}\|\eta_{s}\|\sigma}e^{-C^{\prime\prime}\|\eta_{s}\|a}\|(\mathcal{N}+10)^{2}\psi\|$
$\leq C\|\eta_{s}\|^{-1}M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\ (e^{C\|\eta_{s}\|}-e^{C\|\eta_{s}\|\tau})\|(\mathcal{N}+10)^{2}\psi\|.$ (5.56)
@@
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,5}\psi\|^{2}}\leq CM^{2}N^{-\frac{4}{3}+2\delta}\|(\mathcal{N}+10)^{2}\psi\|.$ (5.57)
@@
$\|I_{4}\psi\|$ $=\|2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{k,\alpha}\psi\|$
$\leq C\int_{0}^{1}\textnormal{d}\tau\sum_{n=1}^{5}\|\eta_{s}\|\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,n}\psi\|^{2}\bigg{)}^{1/2}$
$\leq CMN^{-\frac{2}{3}+\delta}\left((\|\eta_{s}\|^{3}+\|\eta_{s}\|-1)e^{C\|\eta_{s}\|}+1\right)\|(\mathcal{N}+8)^{2}\psi\|$ (5.58)
@@
$=$ $\|\sum_{i=1}^{3}\Big{\{}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s})$
[…]
$\leq$ $C\left((\|\eta_{s}\|+\|\eta_{s}\|^{2})e^{C\|\eta_{s}\|}+(\|\eta_{s}\|^{2}+\|\eta_{s}\|^{4})\right)\sum_{i=1}^{3}\|\partial_{y_{i}}\phi\|+CMN^{-\frac{2}{3}+\delta}e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\sum_{i=1}^{3}\|\partial_{y_{i}}\phi\|$
@@
The second line (6.1b) is be bounded from below by bounds of order $o(1)$ from Theorem 3.1 and Theorem 3.4. Therefore it is sufficient to show that the first line (6.1a) is large.
@@
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\left\{\big{(}\widetilde{\mathbb{H}^{\text{eff}}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\eta_{s},\dot{\eta}_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\right\}$ (6.2)
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\bigg{\{}\big{(}h_{0}-c^{\ast}(h_{y})-c(h_{y})\big{)}W(\eta_{s})\psi+\text{Error}\bigg{\}}$ (6.3)
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\bigg{\{}-W(\eta_{s})\left(c^{\ast}(h_{y})+\langle\eta_{s},h_{y}\rangle+\langle h_{y},\eta_{s}\rangle\right)\psi+h_{0}W(\eta_{s})\psi+\widetilde{\text{Error}}\bigg{\}}$ (6.4)

### v2

@@ 3.2 Effective coherent state
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ \|(\mathcal{N}+3)e^{aB}\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|$
$\leq C\|\eta_{s}\|\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|a}\|(\mathcal{N}+5)^{2}e^{\sigma B}\psi\|$
$\leq C\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma(e^{C\|\eta_{s}\|(1-\sigma)}-1)e^{\tilde{C}\|\eta_{s}\|\sigma}\|(\mathcal{N}+8)^{2}\psi\|$
$\leq C\|\eta_{s}\|^{-1}\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|e^{C\|\eta_{s}\|(1-\tau)}\|(\mathcal{N}+8)^{2}\psi\|$ (5.64)
where we used Cauchy-Schwarz with (5.30), Lemma 5.6 and Lemma 5.8 in the first inequality, Proposition 5.4 and Lemma A.3 in the second inequality. Therefore it follows
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,2}\psi\|^{2}}\leq CM^{2}N^{-\frac{4}{3}+2\delta}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+8)^{2}\psi\|.$
@@
$\|I_{4,3}\psi\|\leq C\|\eta_{s}\|^{-1}MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|(e^{C\|\eta_{s}\|}-e^{C\|\eta_{s}\|\tau})\|(\mathcal{N}+3)\psi\|$ (5.65)
@@
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,3}\psi\|^{2}}\leq CMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+3)\psi\|.$ (5.66)
@@
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \sigma\|\mathcal{N}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\sigma\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)e^{\sigma B}\psi\|$
$\leq C(\|\eta_{s}\|+\|\eta_{s}\|^{2})MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \sigma\|(\mathcal{N}+1)^{2}e^{\sigma B}\psi\|$
$\leq C(\|\eta_{s}\|^{-1}+1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\ \left(e^{C\|\eta_{s}\|}+(1-C\|\eta_{s}\|\tau)e^{C\|\eta_{s}\|\tau}\right)\|(\mathcal{N}+3)^{2}\psi\|$
$\leq C(\|\eta_{s}\|^{-1}+1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\ e^{C\|\eta_{s}\|}\|(\mathcal{N}+3)^{2}\psi\|$ (5.67)
@@
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,4}\psi\|^{2}}$
@@
$\|I_{4,5}\psi\|$ $=\|2i\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\int_{0}^{\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{\Gamma}e^{aB}\psi\|$
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ \|\mathcal{N}e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{\Gamma}e^{aB}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|(1-a)}\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|(\mathcal{N}+3)\mathcal{R}^{\sigma-a}_{\alpha}(k)e^{aB}\psi\|^{2}\bigg{)}^{1/2}$
$\leq C\|\eta_{s}\|M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|(1-a)}(e^{C\|\eta_{s}\|(\sigma-a)}-1)\|(\mathcal{N}+8)^{2}e^{aB}\psi\|$
$\leq C\|\eta_{s}\|M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|}e^{C^{\prime}\|\eta_{s}\|\sigma}e^{-C^{\prime\prime}\|\eta_{s}\|a}\|(\mathcal{N}+10)^{2}\psi\|$
$\leq C\|\eta_{s}\|^{-1}M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\ (e^{C\|\eta_{s}\|}-e^{C\|\eta_{s}\|\tau})\|(\mathcal{N}+10)^{2}\psi\|$ (5.69)
@@
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,5}\psi\|^{2}}\leq CM^{2}N^{-\frac{4}{3}+2\delta}\|(\mathcal{N}+10)^{2}\psi\|.$ (5.70)
@@
$\|I_{4}\psi\|$ $=\|2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{\Gamma}\psi\|$
$\leq C\int_{0}^{1}\textnormal{d}\tau\sum_{n=1}^{5}\|\eta_{s}\|\bigg{(}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,n}\psi\|^{2}\bigg{)}^{1/2}$
$\leq CMN^{-\frac{2}{3}+\delta}\left((\|\eta_{s}\|^{3}+\|\eta_{s}\|-1)e^{C\|\eta_{s}\|}+1\right)\|(\mathcal{N}+8)^{2}\psi\|.$ (5.71)
@@
$=\|\sum_{i=1}^{3}\Big{(}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s})$
$+2i\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})\Big{)}\partial_{y_{i}}\psi\|$
$\leq C\left((\|\eta_{s}\|+\|\eta_{s}\|^{2})e^{C\|\eta_{s}\|}+(\|\eta_{s}\|^{2}+\|\eta_{s}\|^{4})\right)\sum_{i=1}^{3}\|\partial_{y_{i}}\phi\|$
$+CMN^{-\frac{2}{3}+\delta}e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\sum_{i=1}^{3}\|\partial_{y_{i}}\phi\|$ (5.73)
@@
Proofs of properties of $\eta$
Proof of Lemma 5.6.
Recall that by definition it holds
$\|\eta_{s}\|^{2}=\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}|\big{(}\eta_{s}\big{)}_{\alpha}(k)|^{2}=\lambda^{2}\sum_{k\in\Gamma}|\hat{V}(k)|^{2}\underbrace{\sum_{\alpha\in\mathcal{I}_{k}}\bigg{|}n_{\alpha}(k)\frac{\sin\big{(}\epsilon_{\alpha}(k)s/2\big{)}}{\epsilon_{\alpha}(k)/2}\bigg{|}^{2}}_{\eqqcolon S_{k}}$ (5.77)
with $\epsilon_{\alpha}(k)=2k_{F}|k\cdot\omega_{\alpha}|$. We will first approximate $S_{k}$ and then give a upper bound.
$S_{k}$ $=\frac{1}{|k|}\int_{0}^{2\pi}\textnormal{d}\varphi\int_{0}^{\pi/2}\textnormal{d}\theta\frac{\sin(k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\sin\theta+\mathcal{E}$
$=\frac{2\pi}{|k|}\int_{0}^{1}\textnormal{d}u\frac{\sin(k_{\textnormal{F}}|k|su)^{2}}{u}+\mathcal{E}$
$=\frac{\pi}{|k|}\left\{\log(2k_{\textnormal{F}}|k|s)-\text{Ci}(2k_{\textnormal{F}}|k|s)+\gamma\right\}+\mathcal{E}$ (5.79)
$\mathcal{E}_{2}$ $=\left|\int_{p_{\alpha}}\textnormal{d}\sigma\ \frac{\sin(k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}-\sigma(p_{\alpha})\frac{\sin(k_{\textnormal{F}}|k|s\cos\theta_{\alpha})^{2}}{\cos\theta_{\alpha}}\right|$
$\leq\sup_{\hat{\omega}(\theta,\varphi)\in p_{\alpha}}\left|\frac{\textnormal{d}}{\textnormal{d}\theta}\frac{\sin(k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\right|\sup_{(\theta,\varphi)\in p_{\alpha}}\left|\theta-\theta_{\alpha}\right|\sigma(p_{\alpha})$
$\leq C(k_{\textnormal{F}}|k|s)^{2}M^{-\frac{3}{2}}$ (5.80)
$\mathcal{E}_{3}=\sup_{\hat{\omega}(\theta,\varphi)\in p_{\alpha}}\left|\frac{\sin(k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\right|\left|\sigma\Big{(}\bigcup_{\alpha\in\mathcal{I}_{k}}p_{\alpha}\Big{)}-\sigma(S)\right|\leq Ck_{\textnormal{F}}|k|s\left(N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right).$ (5.81)
Thus with the triangle inequality
$\left|S_{k}-\frac{\pi}{|k|}\left\{\log(2k_{\textnormal{F}}|k|s)-\text{Ci}(2k_{\textnormal{F}}|k|s)+\gamma\right\}\right|$
$\leq\frac{C}{|k|}\left\{M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+(k_{\textnormal{F}}|k|s)^{2}M^{-\frac{3}{2}}+k_{\textnormal{F}}|k|s\left(N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right)\right\}.$ (5.82)
We will now give an estimate for $S_{k}$ by approximating for any $m\in\mathbb{R}$
$\frac{\sin(mx)^{2}}{x}=\frac{1}{2}\frac{1-\cos(2mx)}{x}\leq\begin{cases}m^{2}x&\text{if }2mx\in[0,\pi/2],\\ \frac{1}{x}&\text{if }2mx>\pi/2\end{cases}$
where we used $1-x^{2}/2\leq\cos(x)$ for all $x\in[0,\pi/2]$. Therefore
$\frac{2\pi}{|k|}\int_{0}^{1}\textnormal{d}u\frac{\sin(k_{\textnormal{F}}|k|su)^{2}}{u}$
$\leq\frac{2\pi}{|k|}\int_{0}^{1}\textnormal{d}u\ \left\{(k_{\textnormal{F}}|k|s)^{2}u\ \chi(2k_{\textnormal{F}}|k|su\leq\pi/2)+\frac{1}{u}\chi(2k_{\textnormal{F}}|k|su>\pi/2)\right\}$
$\leq\frac{2\pi}{|k|}\int_{0}^{1}\textnormal{d}u\ \left\{(k_{\textnormal{F}}|k|s)^{2}u\ \chi\Big{(}u\leq\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}+\frac{1}{u}\ \chi\Big{(}u>\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}\right\}$
$=\frac{2\pi}{|k|}\left\{\frac{1}{2}(k_{\textnormal{F}}|k|s)^{2}\Big{(}\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}^{2}-\log\Big{(}\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}\right\}\ \chi\Big{(}1>\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}$
$\qquad+\frac{2\pi}{|k|}\frac{1}{2}(k_{\textnormal{F}}|k|s)^{2}\ \chi\Big{(}1\leq\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}$
$=\frac{2\pi}{|k|}\left\{\frac{\pi^{2}}{32}-\log(\pi)+\log(4k_{\textnormal{F}}|k|s)\right\}\ \chi\Big{(}k_{\textnormal{F}}|k|s>\frac{\pi}{4}\Big{)}$
$\qquad+\frac{2\pi}{|k|}\frac{1}{2}(k_{\textnormal{F}}|k|s)^{2}\ \chi\Big{(}k_{\textnormal{F}}|k|s\leq\frac{\pi}{4}\Big{)}$
$\leq\frac{2\pi}{|k|}\min\left\{\frac{1}{2}(k_{\textnormal{F}}|k|s)^{2},\log(4k_{\textnormal{F}}|k|s+1)\right\}.$ (5.83)
Note that $\log(4k_{\textnormal{F}}|k|s+1)\leq\log(4k_{\textnormal{F}}s+1)+\log|k|$ using $\frac{1}{2}x^{2}\leq\log(4x+1)$ for all $x\leq 2$ and $|k|\geq 1$ for all $k\in\mathbb{Z}_{\ast}^{3}$. Thus, it holds
$\|\eta_{s}\|^{2}$ $\leq 2\pi\lambda^{2}\sum_{k\in\mathbb{Z}_{\ast}^{3}}|\hat{V}(k)|^{2}\min\left\{\frac{(k_{\textnormal{F}}s)^{2}}{2}|k|,\frac{\log(4k_{\textnormal{F}}|k|s+1)}{|k|}\right\}$
$\leq 2\pi\lambda^{2}\min\left\{\|\hat{V}(\cdot)^{1/2}\|_{2}^{2}(k_{\textnormal{F}}s)^{2}/2,\|\hat{V}\|_{2}^{2}\left(\log(4k_{\textnormal{F}}s+1)+1\right)\right\}.$ (5.84)
$\|\eta_{s}\|$ $\leq\lambda\min\left\{bk_{\textnormal{F}}s,\log(4k_{\textnormal{F}}s+2)^{a}+a\right\},$ (5.85)
$e^{c_{0}\|\eta_{s}\|}$ $\leq\min\left\{e^{bc_{0}\lambda k_{\textnormal{F}}s},e^{ac_{0}\lambda}(4k_{\textnormal{F}}s+2)^{ac_{0}\lambda}\right\}$ (5.86)
with $a=\sqrt{2\pi}\|\hat{V}\|_{2},b=\sqrt{\pi}\|\hat{V}(\cdot)^{1/2}\|_{2}.$
Proof of Lemma 5.8.
For the inequalities, we observe that for $n\in\mathbb{N}_{0}$ it holds
$\sum_{k\in\Gamma}\||k|^{n}\eta_{s}(k)\|_{l^{2}}$ $\equiv\sum_{k\in\Gamma}|k|^{n}\Big{(}\sum_{\alpha\in\mathcal{I}_{k}}|(\eta_{s})_{\alpha}(k)|^{2}\Big{)}^{1/2}=\lambda\sum_{k\in\Gamma}|k|^{n}|\hat{V}(k)|\left(S_{k}\right)^{1/2}$ (5.87)
with $S_{k}$ from (5.77) Since by assumption $\|(\cdot)^{n}\hat{V}\|_{1}$ is bounded for each $n\in\mathbb{N}$, the first statement follows. The second statement simply follows from the same calculation and recalling that $\langle\eta_{s},|k|^{n}\eta_{s}\rangle\equiv\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}|k|^{n}|\eta_{\alpha}(k)|^{2}$. The third statement follows from Lemma A.4 and
$\|c^{\ast}(|k|^{n}\eta_{s})\psi\|\leq\sum_{k\in\Gamma}\|\sum_{\alpha\in\mathcal{I}_{k}}|k|^{n}\eta_{\alpha}(k)c_{\alpha}^{\ast}(k)\psi\|\leq\sum_{k\in\Gamma}\||k|^{n}\eta_{s}(k)\|_{l^{2}}\|(\mathcal{N}+1)^{1/2}\psi\|.$ (5.88)
@@
The second line (6.1b) is to be bounded from below by bounds of order $o(1)$ from Theorem 3.1 and Theorem 3.6. Therefore it is sufficient to show that the first line (6.1a) is large.
@@
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\left(\big{(}\widetilde{\mathbb{H}^{\text{eff}}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\eta_{s},\dot{\eta}_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\right)$ (6.2)
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\bigg{(}\big{(}h_{0}-c^{\ast}(h_{y})-c(h_{y})\big{)}W(\eta_{s})\psi+\text{Error}\bigg{)}$ (6.3)
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\bigg{(}-W(\eta_{s})\left(c^{\ast}(h_{y})+\langle\eta_{s},h_{y}\rangle+\langle h_{y},\eta_{s}\rangle\right)\psi+h_{0}W(\eta_{s})\psi+\widetilde{\text{Error}}\bigg{)}$ (6.4)